OpenStudy (anonymous):

Which of these produces a finite set of solutions for x? * x2/2 < 35 * x2/2 = 35 * x2/2 ≤ 35 * x2/2 > 35

OpenStudy (anonymous):

Just looking at the symbols in the equations, which one seems different than the others?

OpenStudy (anonymous):

the one with the = sign

OpenStudy (anonymous):

that would be the right answer :)

OpenStudy (anonymous):

how tho? just because it has a different sign than the others?

OpenStudy (anonymous):

well no...i should explain my bad.

OpenStudy (anonymous):

or few can.

OpenStudy (anonymous):

All yours ;) I'm sure you'll do a far superior job.

OpenStudy (anonymous):

joemath i have a doubt here

OpenStudy (anonymous):

i was just exploiting multiple choice questions. its a test taking strat. As for why the one with the = is the answer, its because when we solve for x we obtain:\[\frac{x^2}{2}= 35\Longrightarrow x^2 = 70 \Longrightarrow x = \pm \sqrt{70}\] which gives us only 2 answers, a finite number.

OpenStudy (anonymous):

Hmmm...i think I can see some issues coming up depending if we are talking about real numbers, or natural numbers. If its real numbers, then im correct. If its natural numbers, then there are more correct answers. Is it possible to have more than one correct answer?

OpenStudy (anonymous):

No. I am only allowed to choose one answer on this one. =/

OpenStudy (anonymous):

yeah i am asking about the same thing joe

OpenStudy (anonymous):

ah, since there is only one right answer, i believe what i put is correct. Thanks for catching that Atch, its a poorly worded problem.

OpenStudy (anonymous):

hmmm yeah